Company ABC limited capital structure as December 31st 2020 was as follows.



Amount



Ordinary share 80,000@sh 10 sh 800,000



12.5% preference shares@sh 15 sh 900,000



10% debentures @sh 100 sh 600,000



14% bank loan sh 500,000



Retained earnings sh 450,000



Additional information



The market value of the Equity is sh 15 and firm pays 10% dividend whereas the growth rate was 5% pa.



The market price of preference was sh 25 and the debenture market value was sh 90.



The equity flotation fee per share was sh 1 and corporation tax was 30%.



Required



Compute the cost of each source of fund

1
Expert's answer
2021-11-28T17:16:06-0500

Cost of each source computation

a) Cost of Equity

Cost of Equity=(Dividends Per Share/ Current Market Price Per Share)+Growth Rate of Dividends

Ke= (DO(1 + g)/MPS) + g

10(1+0.05)/15+0.05

10.5/15 + 0.05

=0.7+0.05

=0.75

0.75*100

=75%

b)Cost of preference shares

Kp = Dp/NP

Whereas;

Kp = Cost of Preference Share

Dp = Dividend on preference share

NP = Net proceeds from issue of preference share

Thus,

Kp=(12.5%*25)/25

=0.26

0.26*100

=26%

c)Cost of debentures.

Kd=Interest*(1 – tax rate)+(Par Value - Market Value)1/n/0.5(Par Value + Market Value)

10*(1 - 30/100)+(100 - 90)1/0.5(100 + 90)

(7 + 10)/95

17/95

=0.1790

=17.9%